# Sports-QA: A Large-Scale Video Question Answering Benchmark for Complex and Professional Sports

### v2

@@ 3.4 Dataset Statistics
Table 2: The numbers of QA pairs for different types and different sports. “Desc. / Temp. / Counter.” represents “Descriptive / Temporal / Counterfactual”.
Desc. Temp. Causal Counter. Total
@@
The total number of videos in our dataset is 5,967, distributed across different sports as follows: basketball (879), football (1,030), volleyball (586), gym (505), vault (501), uneven bars (834), floor exercise (770), and balance beam (862). Sports-QA is divided into training, validation, and testing sets (60%20%20%). Specifically, videos of the same sport are randomly assigned to each set to ensure similar distributions of sport types across subsets. Additionally, we have ensured that the distributions of answer classes are similar in different subsets [^4] Please refer to the supplementary material for more statistics..
@@ 4 Auto-Focus Transformer for Sports Video Question Answering
${\rm AFA}(\bm{q}_{j})=\sum_{f\in\mathcal{F}}\alpha_{f}\sum_{i\in\mathcal{D}_{j}^{f}}{\rm softmax}_{i}\left(\bm{q}_{j}^{\mathrm{T}}\bm{k}_{i}/\sqrt{d}\right)\bm{v}_{i},$ (1)
where $\mathcal{F}$ is a set of pre-defined focal lengths. $\mathcal{D}_{j}^{f}=\{i||i-j|\leq f\}$ is the index set of the neighbours of $\bm{q}_{j}$ within focal length $f$. $\alpha_{f}\in(0,1)$ is the dynamic weight of focus $f$, satisfying $\sum_{f}\alpha_{f}=1$ and depended on the question. As shown in Figure 3, $\bm{\alpha}=\{\alpha_{f}\}_{f=1}^{\mathcal{|F|}}$ are obtained by applying linear projection and softmax to the global representation of question. Compared to traditional attention that models global dependencies in any case, our AFA captures only the temporal dependencies of the scale required by the question. This design allows the model to focus solely on crucial information and discard irrelevant parts for better prediction. We denote the output of AFT as $\bm{M}$.
@@ 5.1 Experimental Settings
Pre-processing and Representation. The videos in our dataset are down-sampled to 5 FPS to reduce temporal redundancy. We extract two types of visual features for experiments: appearance feature and motion feature. For each feature type, we obtain both global and local features. Specifically, we detect 10 objects in every two frames of each video using Faster-RCNN [26] (with ResNet-101 [12] as the backbone and a score threshold set to 0.6), pretrained on COCO [21]. These detected objects serve as local appearance features in our work. Additionally, we use the global average pooling of Res4 activation in Faster-RCNN as the global appearance feature. To extract motion features, we employ I3D [3] pretrained on Kinetics [3]. Specifically, for each frame, we combine it with its adjacent 7 frames, creating a clip that we then feed to I3D to extract a 3D feature map. The global motion feature of a frame is obtained by applying global average pooling to the 3D feature. The local motion feature of an object in a frame is obtained by applying RoIAlign [13] (output size $1\times 1$) to the 3D feature with its corresponding bounding box. Note that our method utilizes only the global feature. On the language side, we explore two types of word embeddings: GloVe [24] and pre-trained BERT [6].
@@
Word Emb. Baseline Accuracy F1-score
Random 0.5 0.3
S-A Random 24.6 5.4
GloVe LSTM 44.0 13.5
Transformer 43.7 14.9
BERT LSTM 43.9 13.8
Transformer 43.9 15.8
5.2 Baseline Study
Random Choosing: Answers are randomly selected from the 191 answer classes, resulting in an accuracy of approximately $\frac{1}{191}\approx 0.5$. The F1-score is obtained by running the random test for 200 times and taking the average.
5.3 Benchmark on Sports-QA
We benchmark several VideoQA methods on our Sports-QA dataset, comparing them to CoMem [9], HME [7], HGA [15], MASN [27], HQGA [35], and IGV [20], with HQGA [35] and IGV [20] representing the state of the art. It’s worth noting that we only evaluate methods without large-scale video-text pretraining, excluding pretrained models like MERLOT [41], VIOLET [8], and All-in-one [32]. The exclusion is based on the following reasons: 1) pretrained models require large-scale additional data, making comparisons unfair for methods without pretraining; 2) we aim to highlight the characteristics of our dataset, and the universal knowledge gained by pretrained models could introduce bias in understanding sports scenarios. Table 4 shows the types of visual features used by each model, along with a summary of each method’s insights. For CoMem, HME, and HGA, we use re-implementations provided by [34], and for MASN, HQGA, and IGV, we utilize the official implementations by their respective authors. The baseline model has the same architecture as shown in Figure 3, with AFT replaced with a standard Transformer Encoder.
Model Venue Insight G.A G.M. L.A. L.M.
CoMem [9] CVPR’18 Mem
HME [7] CVPR’19 Mem
HGA [15] AAAI’20 GNN
IGV [20] CVPR’22 IG
HQGA [35] AAAI’22 HL
MASN [27] ACL’21 GNN
AFT (Ours) AFA
@@ 5.3 Benchmark on Sports-QA
5.4 Visualization
Qualitative Results. Figure 4 illustrates predictions from various methods. In some instances, BlindQA successfully guesses the correct answer by leveraging the correlation between the question and the answer in the training set. However, this approach is not foolproof, as the blinded model consistently selects the most correlated answer to the question. In comparison, the VideoQA models struggle to discern subtle differences between semantically or visually similar answer classes. For example, actions like “long pass” and “cross” both involve passing the ball to other players but differ in passing direction and distance, presenting challenges for accurate distinction. Additionally, the compared VideoQA models exhibit limitations in fine-grained motion analysis, such as action counting, as evident in the last example. With its ability to capture various types of temporal dependencies, our AFT outperforms the compared models on these challenging samples.
@@ 5.4 Visualization
Focus Weight. Figure 5 depicts the focus weights of selected questions. Notably, questions seeking information about specific actions (e.g., “win a point”) heavily rely on short-term information (focus length 3). In contrast, questions like “How many times does the right team do ‘block’?” necessitate a stronger emphasis on long-term information. Questions pertaining to the relationships between two actions show a preference for mid-term information, offering insights into the cause and effect of actions. Intuitively, the mid-term focus length can provide more nuanced information on the interactions between actions. This visualization affirms the effectiveness of our model, which dynamically integrates different focus scales based on the diverse temporal dependencies required by various questions [^5] For additional analysis, please refer to the supplementary materials..

### v4

@@ 3.4 Dataset Statistics
Table 3: The numbers of QA pairs for different types and different sports.
Sports Descriptive Temporal Causal Counterfactual Total
@@
[…]
Figure 4: The distributions of answer classes broken down by question types.
The total number of videos in our dataset is 5,967, distributed across different sports as follows: basketball (879), football (1,030), volleyball (586), gym (505), vault (501), uneven bars (834), floor exercise (770), and balance beam (862). Sports-QA is divided into training, validation, and testing sets (60%/20%/20%). Specifically, videos of the same sport are randomly assigned to each set to ensure similar distributions of sport types across subsets. Additionally, we have ensured that the distributions of answer classes are similar in different subsets.
Fig. 4 illustrates the distribution of answer classes broken down by question types in Sports-QA. Note that we only showcase the first 80 classes with more examples. As depicted in the figure, the distribution of answers is long-tailed, and the unbalanced nature adds to the challenge of our dataset. Meanwhile, classes of the same type, such as “yes/no”, are balanced. This further increases the difficulty of our dataset, as the models are required to examine the video for answering the question instead of merely guessing the answer based on the question type.
@@ 4 Auto-Focus Transformer for Sports Video Question Answering
${\rm AFA}(\bm{q}_{j})=\sum_{f\in\mathcal{F}}\alpha_{f}\sum_{i\in\mathcal{D}_{j}^{f}}{\rm softmax}_{i}\left(\frac{\bm{q}_{j}^{\mathrm{T}}\bm{k}_{i}}{\sqrt{d}}\right)\bm{v}_{i},$ (1)
where $\mathcal{F}$ is a set of pre-defined focal lengths. $\mathcal{D}_{j}^{f}=\{i||i-j|\leq f\}$ is the index set of the neighbours of $\bm{q}_{j}$ within focal length $f$. $\alpha_{f}\in(0,1)$ is the dynamic weight of focus $f$, satisfying $\sum_{f}\alpha_{f}=1$ and depended on the question. As shown in Fig. 5, $\bm{\alpha}=\{\alpha_{f}\}_{f=1}^{\mathcal{|F|}}$ are obtained by applying linear projection and softmax to the global representation of question. Compared to traditional attention that models global dependencies in any case, our AFA captures only the temporal dependencies of the scale required by the question. This design allows the model to focus solely on crucial information and discard irrelevant parts for better prediction. We denote the output of AFT as $\bm{M}$.
@@ 5.1 Experimental Settings
Pre-processing and Representation. The videos in our dataset are down-sampled to 5 FPS to reduce temporal redundancy. We extract two types of visual features for experiments: appearance feature and motion feature. For each feature type, we obtain both global and local features. Specifically, we detect 10 objects in every two frames of each video using Faster-RCNN [30] (with ResNet-101 [13] as the backbone and a score threshold set to 0.6), pretrained on COCO [21]. These detected objects serve as local appearance features in our work. Additionally, we use the global average pooling of Res4 activation in Faster-RCNN as the global appearance feature. To extract motion features, we employ I3D [4] pretrained on Kinetics [4]. Specifically, for each frame, we combine it with its adjacent 7 frames, creating a clip that we then feed to I3D to extract a 3D feature map. The global motion feature of a frame is obtained by applying global average pooling to the 3D feature. The local motion feature of an object in a frame is obtained by applying RoIAlign [12] (output size $1\times 1$) to the 3D feature with its corresponding bounding box. Note that our method uses only the global feature. For language, we explore two types of word embeddings: GloVe [28] and pre-trained BERT [6].
@@
Metric Baseline GloVe BERT
Random S-A Random LSTM Transformer LSTM Transformer
Accuracy 0.5 24.6 44 43.7 43.9 43.9
F1-score 0.3 5.4 13.5 14.9 13.8 15.8
@@
[…]
Model Venue Insight G.A G.M. L.A. L.M.
CoMem [10] CVPR’18 Mem
HME [8] CVPR’19 Mem
HGA [14] AAAI’20 GNN
IGV [24] CVPR’22 IG
HQGA [40] AAAI’22 HL
MASN [31] ACL’21 GNN
AFT (Ours) AFA
[…]
5.2 Baseline Study
Random Choosing: Answers are randomly selected from the 191 answer classes, resulting in an accuracy of approximately $\frac{1}{191}\approx 0.5\%$. The F1-score is obtained by running the random test for 200 times and taking the average.
5.3 Benchmark on Sports-QA
We benchmark several VideoQA methods on our Sports-QA dataset, comparing them to CoMem [10], HME [8], HGA [14], MASN [31], HQGA [40], and IGV [24], with HQGA [40] and IGV [24] representing the state of the art. It’s worth noting that we only evaluate methods without large-scale video-text pretraining, excluding pretrained models like MERLOT [48], VIOLET [7], and All-in-one [37]. The exclusion is based on the following reasons: 1) pretrained models require large-scale additional data, making comparisons unfair for methods without pretraining; 2) we aim to highlight the characteristics of our dataset, and the universal knowledge gained by pretrained models could introduce bias in understanding sports scenarios. Table 6 shows the types of visual features used by each model, along with a summary of each method’s insights. For CoMem, HME, and HGA, we use re-implementations provided by [39], and for MASN, HQGA, and IGV, we utilize the official implementations by their respective authors. The baseline model has the same architecture as shown in Fig. 5, with AFT replaced with a standard Transformer Encoder.
@@ 5.3 Benchmark on Sports-QA
5.4 Further Analysis
Generalization Ability across Sports. In this study, our primary objective is to assess the generalization capabilities across different sports domains. Our approach involves an initial pretraining phase on a set of four sports, followed by fine-tuning on a fifth, distinct sport. We meticulously evaluate the performance of this pretraining-fine-tuning methodology in comparison to training a model entirely from scratch. The outcomes, detailed in Table 7, showcase a notable enhancement in accuracy when employing the pretraining strategy on diverse sports. These findings suggest that the model acquires a nuanced understanding of common semantic features during the initial pretraining. Moreover, the model demonstrates its ability to effectively generalize this acquired knowledge to previously unseen sports during the subsequent fine-tuning phase.
Table 7: The table presents comparisons (accuracy %) between models trained from scratch for individual sports and those fine-tuned using data from other sports, along with the corresponding improvements (Imp.).
Pretraining Basketball Football Volleyball Gym FineG.
57.1 55.0 56.4 70.2 48.5
58.3 56.5 57.7 71.7 50.3
Imp. +1.2 +1.5 +1.3 +1.5 +1.8
[…]
Table 8: The accuracy (%) of various focus lengths and their combinations. Note that the combination of the same focal length implies the use of only one focus length.
1 3 5 7 9 11 80
1 58.2 / / 58.9 59.0 58.7 58.5
3 / 58.7 / 59.0 59.1 58.6 58.9
5 / / 58.6 58.8 58.9 58.7 58.8
7 / / / 58.7 / / 58.9
9 / / / / 58.8 / 59.0
11 / / / / / 58.4 58.6
80 / / / / / / 58.0
[…]
Impact of Focus Length. In this experiment, we investigate the impact of different focus lengths and their combinations. For short-term dependency, we consider lengths of 1, 3, and 5. For the mid-term, we explore lengths of 7, 9, and 11. Regarding long-term (global) dependency, we set the focus length to the video length, which is 80 seconds. The results are presented in Table 8. From the results, it is evident that the model with global attention performs less effectively compared to those with local attention. However, upon applying our auto-focus attention mechanism, there is a significant improvement in performance. Notably, the combination of focus lengths 3 and 9 achieves the highest accuracy. In our final model, we incorporate the global focus, resulting in the best accuracy at 59.2%.
5.5 Visualization
Qualitative Results. Fig. 6 illustrates predictions from various methods. In some instances, BlindQA successfully guesses the correct answer by leveraging the correlation between the question and the answer in the training set. However, this approach is not foolproof, as the blinded model consistently selects the most correlated answer to the question. In comparison, the VideoQA models struggle to discern subtle differences between semantically or visually similar answer classes. For example, actions like “long pass” and “cross” both involve passing the ball to other players but differ in passing direction and distance, presenting challenges for accurate distinction. Additionally, the compared VideoQA models exhibit limitations in fine-grained motion analysis, such as action counting, as evident in the last example. With its ability to capture various types of temporal dependencies, our AFT outperforms the compared models on these challenging samples.
@@ 5.5 Visualization
Focus Weight. Fig. 7 depicts the focus weights of selected questions. Notably, questions seeking information about specific actions (e.g., “win a point”) heavily rely on short-term information (focus length 3). In contrast, questions like “How many times does the right team do ‘block’?” necessitate a stronger emphasis on long-term information. Questions pertaining to the relationships between two actions show a preference for mid-term information, offering insights into the cause and effect of actions. Intuitively, the mid-term focus length can provide more nuanced information on the interactions between actions. This visualization affirms the effectiveness of our model, which dynamically integrates different focus scales based on the diverse temporal dependencies required by various questions.
@@ 6 Conclusion
7 Limitation and Future Work
In our study, we introduce the innovative Auto-Focus Transformer (AFT) designed to capture intricate dependencies across frames within diverse temporal windows, catering to a range of question types. Unlike the conventional Transformer model, the AFT demands increased computational resources, scaling linearly with the number of focuses. Although we have optimized the implementation for high parallelization, it may pose challenges on devices with constrained memory capacities. Looking ahead, our research will focus on devising a novel architectural framework that retains the distinctive features of AFT while ensuring computational efficiency, addressing the limitations associated with memory constraints.
8 Ethical Consideration
In developing our Sports-QA dataset, sourced from MultiSports and FineGym, we prioritized ethical considerations, adhering to data privacy and usage standards while ensuring compliance with the terms of the original datasets. To address potential biases, we actively promoted diversity in sports categories and specific sports actions, aiming for a balanced representation. Mindful of the environmental impact, we optimized our data processing to minimize resource consumption. Acknowledging the potential for misuse, we emphasize the dataset’s intended use in positive sports analytics applications, such as performance analysis. Our commitment to transparency and reproducibility is evident in our methodology and dataset accessibility, ensuring that our contribution to sports analytics and computer vision is both responsible and beneficial.
List of Crucial Actions
The selected crucial actions are listed in Table 9.
Table 9: Crucial actions selected for Sports-QA, which are classified into offensive actions and defensive actions.
Sports Action
Offensive Defensive
Volleyball serve first pass
dink defend
spike protect
second attack adjust
no offensive attack save
Football shoot diving
cross tackle
through pass steal
long pass block
aerial duels
Basketball jump ball save
free throw pass steal
2-point shot dribble steal
3-point shot defensive rebound
offensive rebound
[…]
Question Templates for Sports-QA
Gymnastics
Descriptive Question
  • What is the video about?
  • How many SOME-ACTION do the players perform?
  • How many types of SOME-ACTION do the players perform?
  • How many times do the players perform SOME-ACTION?
  • How many players perform the i-th SOME-ACTION?
  • Do the players perform SOME-ACTION?
Temporal Question
  • Do the players perform SOME-ACTION before/after the i-th SOME-ACTION?
  • What do the players perform before/while/after performing the i-th SOME-ACTION?
  • How many times do the player do SOME-ACTION before/after doing the i-th SOME-ACTION?
Ball Games
Descriptive Question
  • What is the video about?
  • Does one team score in the video?
  • Which team scores in the video?
  • How many times does SOME-TEAM do SOME-ACTION?
  • How many times does SOME-TEAM do SOME-ACTION successfully?
  • Does SOME-TEAM do their i-th SOME-ACTION successfully?
Temporal Question
  • What does SOME-TEAM do before/after their i-th SOME-ACTION?
  • What does SOME-TEAM do before/after the other team does their i-th SOME-ACTION?
Causal Question
  • Why does SOME-TEAM do the i-th SOME-ACTION?
  • What is the effect of the i-th SOME-ACTION of SOME-TEAM?
  • How does SOME-TEAM succeed in doing the i-th something?
Counterfactual Question
  • Would SOME-TEAM succeed in doing the i-th SOME-ACTION if the other team did not do SOME-ACTION?
Declarations
Data and Code Availability
The proposed dataset, Sports-QA, and the code of the proposed method are available at https://github.com/HopLee6/Sports-QA.